# Riemannian and Lorentzian Calderón problem under Magnetic Perturbation

### v1

@@ 3.1. Preliminaries
We will be working on admissible Lorentzian manifolds, the definition is a slight modification of [19, Definition 1.1].
@@
Let $n\geq 3$. Let $(M,g)$ be a smooth connected $n$-dimensional Lorentzian manifold with non-empty boundary; thus, $g$ has signature (,+,,+)(-,+$\dots$,+)( - , + , … , + ). We call $(M,g)$ admissible if
@@
We briefly introduce concepts in microlocal analysis, for details we refer to [21, 14, 22, 23, 13, 18]. Given a smooth submanifold $K$, we denote $N^{*}K$ its conormal bundle. A distribution $u$ is called a conormal distribution if locally it can be written as an oscillatory integral of the form
@@
where $\phi$ is the phase function locally parameterizing $N^{*}K$, and $a$ is a smooth symbol, see [21]. More generally, when $N^{*}K$ is replaced with some Lagrangian submanifold of $T^{*}M$, the distribution is called a Lagrangian distribution. An operator mapping functions on $X$ to functions on $Y$ is called a Fourier Integral Operator (FIO) if the corresponding Schwartz kernel is a Lagrangian submanifold of $T^{*}Y\times T^{*}X$ with respect to the twisted symplectic form. In this case the Lagrangian submanifold is also referred to as a canonical relation. When $X=Y$ and the Lagrangian submanifold is the diagonal of $T^{*}X\times T^{*}X$, we call the operator a pseudodifferential operator. The wavefront set of any Lagrangian distribution is a subset of the Lagrangian submanifold, see [22, 23].
For any operator $P$ acting on distributions on a Lorentzian or Riemannian manifold $(M,g)$, there exists a natural way to turn it into an operator acting on half-density valued distributions (for definition of half-density, see [18]). For any half-density valued distribution $\omega$, define
$P\omega=|g|^{\frac{1}{4}}P(|g|^{-\frac{1}{4}}\omega).$
Here $|g|^{\frac{1}{4}}=|g|^{\frac{1}{4}}_{x}|dx|^{\frac{1}{2}}$ where $|g|_{x}$ is the determinant of $g$ in local coordinate $x$, and $|dx|^{\frac{1}{2}}$ is the standard half-density in coordinate $x$. Note that this conjugation obviously depends on the metric, one can also conjugate $P$ with other half-densities to avoid this. The benefit of turning $P$ into an operator on half-density valued distribution is that if $P$ is an FIO, then the principal symbol of $P$, denoted by $\sigma(P)$, can be invariantly defined modulo lower order terms as a half-density on the Lagrangian submanifold (for simplicity we ignore the Maslov bundle), see [21] for details.
When the Dirichlet data is a conormal distribution corresponding to some conormal bundle $N^{*}K$, then the solution is a Lagrangian distribution corresponding to the flowout Lagrangian of $N^{*}K$. The flowout Lagrangian refers to the union of the null-bicharacteristics originated from $N^{*}K\cap\mathcal{H}$ where $\mathcal{H}$ is the hyperbolic set (for definition of hyperbolic set and null-bicharacteristic, we refer to [22, Section 24]). We also refer to [40] for an optics construction of the solution.
The DN map $\Lambda_{g,A,q}$ has been proved to contain both pseudodifferential part and Fourier Integral Operator part, see for example [40] and [47]. The corresponding canonical relation for the FIO part is given by the powers of lens relation. Denote $\mathcal{H}$ the hyperbolic set and $\pi:T^{*}M|_{\partial M}\to T^{*}\partial M$. Lens relation on a non-trapping admissible Lorentzian manifold is the map from $\mathcal{H}$ to itself defined in the following way: for any $(x^{\prime},\xi^{\prime})\in\mathcal{H}$, let $\Gamma$ be the unique future pointing null-bicharacteristic from $(x,\xi^{\prime})$, and $(y,\eta)$ is the exit point of $\Gamma$, then $L(x^{\prime},\xi^{\prime})=\pi(y,\eta)$. The projection of $\Gamma$ on $M$ is a null-geodesic, and we will refer to this null-geodesic as the unique geodesic corresponding to $(y^{\prime},\eta^{\prime};x^{\prime},\xi^{\prime})$. For details of lens relation and DN map on a Lorentzian manifold, see for example [40] and [47]. Let $U,V\subset\partial M$ be disjoint, denote $\Lambda^{U,V}_{g,A,q}$ the DN map restricted to $V\times U$, that is the Dirichlet data $f$ in (1.6) is supported in $U$ and
@@
In [47], the principal symbol of $\Lambda^{U,V}_{g,A,q}$ is computed.
Proposition 3.2.
@@
  • the sign is $+$/-- if $(x_{0}^{\prime},\xi^{0\prime})$ is a past/future pointing covector.
@@
$\mathcal{L}_{H_{p}}\sigma(\widetilde{u})=i\sigma(\widetilde{v}).$
For details see [14].
@@ 3.2. Perturbation of The Electromagnetic Potential
Let $L(x^{\prime},\xi^{\prime})=(y^{\prime},\eta^{\prime})$, then they are connected by a unique null-bicharacteristic, whose projection $\gamma$ is a null-geodesic. Note that perturbing $A$ does not change the trajectory of the null-geodesic, but only adds spin to it. In particular, by Proposition 3.2 this means for any $A\in\mathcal{B}_{\delta}(\mathcal{A})$,
@@
On the other hand, suppose $x=\gamma(t_{0})$ for some $t_{0}$, and let $U$ be some neighborhood of $x$, then $\gamma(t)\in U$ for $t\in[t_{0}-\varepsilon,t_{0}+\varepsilon]$ and some $\varepsilon>0$. Consider some $\phi\in C^{\infty}_{c}(\mathbb{R})$ such that supp ($\hbox{supp }(\phi)\subset(-\varepsilon,\varepsilon)$ and $\int\phi=1$. Let $A^{\prime}$ be such that
@@
then it is clear that it holds in a sufficiently small neighborhood of $A$ in $\widetilde{B}_{\delta}(\mathcal{A};U)$ with respect to $C^{k}$ topology. On the other hand, to show density, if $A$ satisfies (3.0), same argument shows that there exists some $B$ arbitrarily close to $A$ such that supp ($\hbox{supp }(B-A)\subset U$ and
@@
It is easy to see from the proof that when constructing the perturbation, we don’t want $A-\mathcal{A}$ to be an exact 1-form supported in the interior, otherwise the integral along the geodesic would still give 0. Another way to see this is that if $A-\mathcal{A}=d\phi$ for some compactly supported $\phi$ in the interior, then the DN map is the same for $A$ and $\mathcal{A}$ by the gauge equivalence (see [40, Lemma 2.2]). A change by closed form, on the other hand may change the DN map, and one can use them to detect trajectory. This aligns with the Aharonov-Bohm effect in physics (see [1, 2]). We essentially need to perturb $A$ outside of the gauge equivalence class, so that DN map would change and give useful information.
@@ 3.3. Perturbation of The Metric
where $q_{g_{\epsilon}}=q_{g}+\epsilon q^{1}+O(\epsilon^{2})$, $\hbox{tr}\,_{g}(h)=\hbox{tr}\,(g^{-1}h)$, and $h_{g}^{jk}$ is the $jk$-entry of the matrix $g^{-1}hg^{-1}$. Note that $q^{1}$ depends on $g$ and $h$, and since the potential depends locally on the metric, supp ($\hbox{supp }(q^{1})\subset\hbox{supp }(h)$. Let $P_{g}=|g|^{\frac{1}{4}}\Box_{g}|g|^{-\frac{1}{4}}$, $\widetilde{u}=u|g|^{\frac{1}{4}}$ and $\widetilde{v}=v|g|^{\frac{1}{4}}$, the equation for the half-densities is
@@
Let $L(x^{\prime},\xi^{\prime})=(y^{\prime},\eta^{\prime})$ and $\gamma$ the unique null-geodesic corresponds to $(y^{\prime},\eta^{\prime};x^{\prime},\xi^{\prime})$. Let $K\subset\partial M$ be such that the conormal bundle $N^{*}K$ contains $(x^{\prime},\xi^{\prime})$. If $f=u|_{\partial M}$ is a conormal distribution with respect to $N^{*}K$, then $u$ is a Lagrangian distribution with canonical relation being the flowout Lagrangian from the hyperbolic part of $N^{*}K$ along the Hamiltonian vector fields (see for example [20], or the optics construction in [40]). As a result, $v$ is a Lagrangian distribution on the same flowout Lagrangian (one can show this by a symbolic construction as in [32], see also [47]). In particular, the principal symbol satisfies the following transport equation
$\mathcal{L}_{H_{p}}\sigma(\widetilde{v})=ih_{g}^{jk}\xi_{j}\xi_{k}\sigma(\widetilde{u})$
where $H_{p}$ is the Hamiltonian vector field with respect to $g$ and $\mathcal{L}_{H_{p}}$ is the Lie derivative along the flowout Lagrangian (see for example [14]). Let $w_{1},\dots,w_{n}$ be linearly independent vector fields on the flowout Lagrangian obtained via pushforward along the Hamiltonian flow $\varphi_{t}$, that is
@@
for some generic compactly supported $\phi$, in any semi-geodesic normal coordinate (see for example [40, Lemma 2.3] or [34])

### v2

@@ 3.1. Preliminaries
We will be working on admissible Lorentzian manifolds, the definition is a slight modification of [HU19, Definition 1.1].
@@
Let $n\geq 3$. Let $(M,g)$ be a smooth connected $n$-dimensional Lorentzian manifold with non-empty boundary; thus, $g$ has signature (,+,,+)(-$\dots,+$). We call $(M,g)$ admissible if
@@
We briefly introduce concepts in microlocal analysis, for details we refer to [Hor71, DH72, Hor85III, Hor85IV, Dui96, Hin25]. Given a smooth submanifold $K$, we denote $N^{*}K$ its conormal bundle. A distribution $u$ is called a conormal distribution if locally it can be written as an oscillatory integral of the form
$u(x)=\int e^{i\phi(x,\theta)}a(x,\theta)d\theta$
where $\phi$ is the phase function locally parameterizing $N^{*}K$, and $a$ is a smooth symbol, see [Hor71]. More generally, when $N^{*}K$ is replaced with some Lagrangian submanifold of $T^{*}M$, the distribution is called a Lagrangian distribution. An operator mapping functions on $X$ to functions on $Y$ is called a Fourier Integral Operator (FIO) if the corresponding Schwartz kernel is a Lagrangian submanifold of $T^{*}Y\times T^{*}X$ with respect to the twisted symplectic form. In this case the Lagrangian submanifold is also referred to as a canonical relation. When $X=Y$ and the Lagrangian submanifold is the diagonal of $T^{*}X\times T^{*}X$, we call the operator a pseudodifferential operator. The wavefront set of any Lagrangian distribution is a subset of the Lagrangian submanifold, see [Hor85III, Hor85IV].
For any operator $P$ acting on distributions on a Lorentzian or Riemannian manifold $(M,g)$, there exists a natural way to turn it into an operator acting on half-density valued distributions (for definition of half-density, see [Hin25]). For any half-density valued distribution $\omega$, define
$\widetilde{P}\omega=|g|^{\frac{1}{4}}P(|g|^{-\frac{1}{4}}\omega).$
@@
When the Dirichlet data is a conormal distribution corresponding to some conormal bundle $N^{*}K$, then the solution is a Lagrangian distribution corresponding to the flowout Lagrangian of $N^{*}K$. The flowout Lagrangian refers to the union of the null-bicharacteristics originated from $N^{*}K\cap\mathcal{H}$ where $\mathcal{H}$ is the hyperbolic set (for definition of hyperbolic set and null-bicharacteristic, we refer to [Hor85III, Section 24]). We also refer to [SY18] for an optics construction of the solution.
The DN map $\Lambda_{g,A,q}$ has been proved to contain both pseudodifferential part and Fourier Integral Operator part, see for example [SY18] and [YZ25]. The corresponding canonical relation for the FIO part is given by the powers of lens relation. Denote $\mathcal{H}$ the hyperbolic set and $\pi:T^{*}M|_{\partial M}\to T^{*}\partial M$. We can define the lens relation $L$ on a non-trapping admissible Lorentzian manifold as follows.
Definition 3.2.
Let $(M,g)$ be a non-trapping admissible Lorentzian manifold. For any $(x^{\prime},\xi^{\prime})\in\mathcal{H}$, there exists a unique inward pointing lightlike direction $\xi$ such that $\pi(x^{\prime},\xi)=(x^{\prime},\xi^{\prime})$. Let $\Gamma$ be the unique future pointing null-bicharacteristic from $(x^{\prime},\xi)$, and $(y^{\prime},\eta)$ is the exit point of $\Gamma$. The lens relation is the map
$L:\mathcal{H}\to\mathcal{H},\quad(x^{\prime},\xi^{\prime})\to\pi(y^{\prime},\xi^{\prime}).$
The projection of $\Gamma$ on $M$ is a null-geodesic, and we will refer to this null-geodesic as the unique geodesic corresponding to $(y^{\prime},\eta^{\prime};x^{\prime},\xi^{\prime})$.
For more details of lens relation and DN map on a Lorentzian manifold, see for example [SY18] and [YZ25]. Let $U,V\subset\partial M$ be disjoint, denote $\Lambda^{U,V}_{g,A,q}$ the DN map restricted to $V\times U$, that is the Dirichlet data $f$ in (1.6) is supported in $U$ and
@@
In [YZ25], the principal symbol of $\Lambda^{U,V}_{g,A,q}$ is computed.
Proposition 3.3 ([YZ25, Proposition 3.5]).
@@
  • the sign is ++/- if $(x_{0}^{\prime},\xi^{0\prime})$ is a past/future pointing covector.
@@
[…]
where $H_{p}$ is the Hamiltonian vector field with respect to the principal symbol of $P_{g}$ and $\mathcal{L}_{H_{p}}$ is the Lie derivative along the flowout Lagrangian. We refer to [DH72] for more details.
@@ 3.2. Perturbation of The Electromagnetic Potential
By [SY18, Theorem 4.3], the lens relation $L$ can be recovered from the DN map. Let $L(x^{\prime},\xi^{\prime})=(y^{\prime},\eta^{\prime})$, then they are connected by a unique null-bicharacteristic, whose projection $\gamma$ is a null-geodesic. Note that perturbing $A$ does not change the trajectory of the null-geodesic, but only adds spin to it. In particular, by Proposition 3.3 this means for any $A\in\mathcal{B}_{\delta}(\mathcal{A})$,
@@
On the other hand, suppose $x=\gamma(t_{0})$ for some $t_{0}$, and let $U$ be some neighborhood of $x$, then $\gamma(t)\in U$ for $t\in[t_{0}-\varepsilon,t_{0}+\varepsilon]$ and some $\varepsilon>0$. Consider some $\phi\in C^{\infty}_{c}(\mathbb{R})$ such that $\hbox{supp }(\phi)\subset(-\varepsilon,\varepsilon)$ and $\int\phi=1$. Let $A^{\prime}$ be such that
@@
then it is clear that it holds in a sufficiently small neighborhood of $A$ in $\widetilde{B}_{\delta}(\mathcal{A};U)$ with respect to $C^{k}$ topology. On the other hand, to show density, if $A$ satisfies (3.0), same argument shows that there exists some $B$ arbitrarily close to $A$ such that $\hbox{supp }(B-A)\subset U$ and
@@
It is easy to see from the proof that when constructing the perturbation, we don’t want $A-\mathcal{A}$ to be an exact 1-form supported in the interior, otherwise the integral along the geodesic would still give 0. Another way to see this is that if $A-\mathcal{A}=d\phi$ for some compactly supported $\phi$ in the interior, then the DN map is the same for $A$ and $\mathcal{A}$ by the gauge equivalence (see [SY18, Lemma 2.2]). A change by closed form, on the other hand may change the DN map, and one can use them to detect trajectory. This aligns with the Aharonov-Bohm effect in physics (see [AB59, AB61]) , which roughly speaking states that the particle may experience a phase shift caused by the electromagnetic potential even when the induced electromagnetic field is 0. Mathematically, $A$ provides a connection on the principal $U(1)$-bundle, and $F=dA$ is the curvature of the connection. The Aharonov-Bohm effect is simply saying that the holonomy depends not just on the curvature, but also the topology of the path.
@@ 3.3. Perturbation of The Metric
where $q_{g_{\epsilon}}=q_{g}+\epsilon q^{1}+O(\epsilon^{2})$, $\hbox{tr}\,_{g}(h)=\hbox{tr}\,(g^{-1}h)$, and $h_{g}^{jk}$ is the $jk$-entry of the matrix $g^{-1}hg^{-1}$. Note that $q^{1}$ depends on $g$ and $h$, and since the potential depends locally on the metric, $\hbox{supp }(q^{1})\subset\hbox{supp }(h)$. Let $P_{g}=|g|^{\frac{1}{4}}\Box_{g}|g|^{-\frac{1}{4}}$, $\widetilde{u}=u|g|^{\frac{1}{4}}$ and $\widetilde{v}=v|g|^{\frac{1}{4}}$, the equation for the half-densities is
@@
Let $L(x^{\prime},\xi^{\prime})=(y^{\prime},\eta^{\prime})$ and $\gamma$ be the unique null-geodesic corresponds to $(y^{\prime},\eta^{\prime};x^{\prime},\xi^{\prime})$. Let $K\subset\partial M$ be such that the conormal bundle $N^{*}K$ contains $(x^{\prime},\xi^{\prime})$. If $f=u|_{\partial M}$ is a conormal distribution with respect to $N^{*}K$, then $u$ is a Lagrangian distribution with canonical relation being the flowout Lagrangian from the hyperbolic part of $N^{*}K$ along the Hamiltonian vector fields (see for example [HUZ22], or the optics construction in [SY18]). As a result, $v$ is a Lagrangian distribution on the same flowout Lagrangian (one can show this by a symbolic construction as in [MU79], see also [YZ25]). In particular, the principal symbol satisfies the following transport equation
[…]
where recall that $H_{p}$ is the Hamiltonian vector field with respect to $g$ and $\mathcal{L}_{H_{p}}$ is the Lie derivative along the flowout Lagrangian. Let $w_{1},\dots,w_{n}$ be linearly independent vector fields on the flowout Lagrangian obtained via pushforward along the Hamiltonian flow $\varphi_{t}$, that is
@@
for some generic compactly supported $\phi$, in any semi-geodesic normal coordinate (see for example [SY18, Lemma 2.3] or [Pet69])
@@
Declarations
The author states that there is no conflict of interest, and there is no associated data used in the paper.